At what positive point do x=y/2+6 and x=y^2 intersect?
Answer by mananth(16946)

x=y/2+6 and x=y^2
Graphically plot points for both graphs and locate the intesection
x=y/2+6
x= 0 , y= y=-12 (0,-12)
x=6 , y=0(6,0)
and x=y^2
x=0, y=0 (0,0)
x=1,y=1 (1,1)
x=4 , y= +/-2 (4,-2) (4,2)
x=9, y=+/-3 (9,3) (9,-3)

The intersection in the first quadrant is the solution. (7.34,2.71)